OpenStudy (killershark737):

A cylinder and a cone have the same diameter: 8 inches. The height of the cylinder is 5 inches. The height of the cone is 15 inches. Use π = 3.14. What is the relationship between the volume of this cylinder and this cone? Explain your answer by determining the volume of each and comparing them. Show all your work.

OpenStudy (dumbcow):

cone is 1/3 volume of similar cylinder since they have same diameter, the bases are the same cone has height 3 times as high as cylinder, this makes volumes equal. \[V_{cone} =\frac{1}{3} \pi r^2 (3h) = \pi r^2 h = V_{cylinder}\] if h = 5